# EEGEncoder: Advancing BCI with Transformer-Based Motor Imagery Classification

### v1

@@ Abstract
Brain-computer interfaces (BCIs) harness electroencephalographic signals for direct neural control of devices, offering a significant benefit for individuals with motor impairments. Traditional machine learning methods for EEG-based motor imagery (MI) classification encounter challenges such as manual feature extraction and susceptibility to noise. This paper introduces EEGEncoder, a deep learning framework that employs transformer models to surmount these limitations. Our innovative multi-scale fusion architecture captures both immediate and extended temporal features, thereby enhancing MI task classification precision. EEGEncoder’s key innovations include the inaugural application of transformers in MI-EEG signal classification, a mixup data augmentation strategy for bolstered generalization, and a multi-task learning approach for refined predictive accuracy. When tested on the BCI Competition IV dataset 2a, our model established a new benchmark with its state-of-the-art performance. EEGEncoder signifies a substantial advancement in BCI technology, offering a robust, efficient, and effective tool for transforming thought into action, with the potential to significantly enhance the quality of life for those dependent on BCIs.
@@ 3 Method
The overall architecture of EEGEncoder is depicted in Figure 3; the subsequent sections will provide a detailed description of the structure and function of each module.
3.1 Patch Projector for EEG Signal Preprocessing
The Patch Projector module within our EEG-based deep learning framework is designed to preprocess Motor Imagery EEG data, preparing it for intricate analysis by subsequent Transformer and Temporal Convolutional Network (TCN) layers. This module adeptly reshapes high-dimensional EEG sequences, characterized by a temporal resolution of 1125 and 22 channels, into a format that is conducive to convolutional processing. The main purpose of this process is to reduce the length of the sequence by passing continuous EEG signals through simple convolutional layers and average pooling layers.
@@ 3.1 Patch Projector for EEG Signal Preprocessing
The core architecture of the Patch Projector, as illustrated in Figure 1, comprises three convolutional layers. The first convolutional layer is designed to initiate the feature extraction process without the application of an activation function. In contrast, the second and third convolutional layers are each followed by a batch normalization (BN) layer and an exponential linear unit (ELU) activation layer to stabilize the learning process and introduce non-linear dynamics into the model. The ELU activation function is defined as:
@@
Figure 1: Architecture of the Patch Projector. The figure provides a detailed schematic of the Patch Projector’s architecture. It includes three convolutional layers, with the second and third layers each followed by a batch normalization (BN) layer and an ELU activation layer. Additionally, two average pooling layers and two dropout layers are incorporated to foster model generalization. Specific parameters, such as the kernel size and stride for the convolutional layers, and the kernel size for the average pooling layers, are also depicted. For example, "Conv 1x16, (64,1)" signifies a convolutional layer transitioning from an input channel depth of 1 to an output channel depth of 16, with a stride of 64 along the width and 1 along the height of the input feature map.
@@ 3.2 Stabilizing the Transformer Layer
In the subsequent modules, we employ a modified Transformer layer, which has been adapted with recent technological advancements to enhance training stability and model efficacy. Here, we detail the specific alterations applied to the Transformer architecture.
@@
def transformer_block_pre_norm(x):     x = norm(x) + self_attention(x)     x = norm(x) + feed_forward(x)     return x
@@
Our approach also incorporates RMSNorm, or Root Mean Square Layer Normalization, as the normalization function. RMSNorm diverges from traditional Layer Normalization by normalizing solely the standard deviation of activations, not the mean and standard deviation. It achieves this by dividing the activations by their root mean square, which maintains gradient scale and facilitates the training of deep networks.
@@
To further enhance our model, we have substituted the conventional ReLU activation with the Sandwiched Gated Linear Unit (SwiGLU). This function introduces an additional non-linearity between two linear transformations:
$\text{SwiGLU}(x)=x\odot\sigma(W_{2}\cdot\text{norm}(W_{1}\cdot x))$ (3)
In the equation, $x$ is the input, $W_{1}$ and $W_{2}$ are weight matrices, $\sigma$ is the sigmoid function, and norm is the non-linear activation. The element-wise product with the sigmoid’s output enables selective feature amplification or damping. SwiGLU’s principal advantages are:
@@
We acknowledge that the concepts of pre-normalization, RMSNorm, and SwiGLU are not our novel contributions . These methods have been previously introduced in works such as Llama and others [24, 25].
@@ 3.3 Dual-Stream Temporal-Spatial Block
The Dual-Stream Temporal-Spatial Block (DTDS Block) presents an architecture specifically designed for the analysis of electroencephalogram (EEG) data during Motor Imagery (MI) tasks. This architecture integrates Temporal Convolutional Networks (TCNs) with stable Transformer modules, capitalizing on their complementary strengths to capture the temporal and spatial characteristics inherent in EEG signals.
@@
The DTDS Block is engineered to leverage the TCN’s proficiency in local feature extraction and the Transformer’s capacity for global context comprehension, thus aiming to provide a comprehensive analysis of EEG data. This dual-stream approach is anticipated to improve the model’s ability to identify patterns relevant to MI tasks by enhancing its analytical complexity.
@@
Figure 2: Architecture of the DTDS Block. The DTDS Block integrates a TCN for local temporal feature extraction with a Transformer for global spatial context analysis, enabling a detailed examination of EEG signals for MI classification tasks.
@@ 3.5 Training Configuration
4 Results
In this section, we provide a detailed evaluation of the EEGEncoder model, demonstrating its classification capabilities on the BCI Competition IV 2a dataset. We compare the performance of our model with various established models to underscore its effectiveness in decoding the complex patterns inherent in EEG signals for motor imagery tasks. The subsequent subsections elaborate on the model’s performance metrics, a comparative analysis with other models, and discuss the significance of these results for the progression of brain-computer interface technologies.
4.1 Datasets
In our study, we primarily utilized the BCI Competition IV dataset 2a (BCI-2A) for training and evaluating the EEGEncoder model. The BCI-2A dataset comprises recordings from nine subjects across two sessions, with each subject performing 288 motor imagery trials. The trials were recorded using 22 EEG electrodes at a sampling rate of 250 Hz, with each trial lasting 4 seconds, resulting in 1125 time points per trial. The data underwent bandpass filtering between 0.5 and 100 Hz and a notch filter at 50 Hz to reduce power line interference.
In contrast, the BCI Competition IV dataset 2b (BCI-2B) focuses on a binary classification problem involving left and right hand movements. This dataset includes three bipolar EEG recordings (C3, Cz, and C4) at a sampling rate of 250 Hz from nine subjects, with five sessions per subject. The initial two sessions were conducted without feedback, consisting of 120 trials each, while the subsequent three sessions included smiley feedback, with 160 trials each.
@@ 4.3 Ablation Study
Here, we present a selection of key experiments that were instrumental in assessing the impact of specific modifications. These experiments included removing the transformer component from the DTDS block, varying the number of transformer layers, adjusting the quantity of DTDS branches within the EEGEncoder, and comparing the performance of our modified stable transformer against the Vanilla Transformer. To ensure the statistical significance of our results, we averaged the outcomes across five iterations, each initialized with a different random seed. The summarized results are displayed in Table 3.
@@
Transformer-2 Layer 74.05 65.4
Transformer-8 Layer 73.78 65.1

### v2

@@ Abstract
Brain-computer interfaces (BCIs) harness electroencephalographic signals for direct neural control of devices, offering a significant benefit for individuals with motor impairments. Traditional machine learning methods for EEG-based motor imagery (MI) classification encounter challenges such as manual feature extraction and susceptibility to noise.This paper introduces EEGEncoder, a deep learning framework that employs modified transformers and TCNs to surmount these limitations. We innovatively propose a fusion architecture, namely Dual-Stream Temporal-Spatial Block (DSTS), to capture temporal and spatial features, improving the accuracy of Motor Imagery classification task. Additionally, we use multiple parallel structures to enhance the model’s performance. When tested on the BCI Competition IV-2a dataset, our model results outperform current state-of-the-art techniques.
Keywords motor imagery  $\cdot$ classification  $\cdot$ electroencephalography (EEG)  $\cdot$ tansformer  $\cdot$ TCN  $\cdot$ Dual-Stream Temporal-Spatial Block (DSTS)
@@ 3 Method
The overall architecture of EEGEncoder is depicted in Figure 1. The proposed EEGEncoder model primarily consists of a Downsampling Projector and multiple parallel Dual-Stream Temporal-Spatial(DSTS) blocks, where the DSTS Block is formed by combining TCN and stable transformer. The Downsampling Projector preprocesses the MI-EEG signals through multiple layers of convolution. DSTS block integrates TCN and stable transformer to extract the temporal and spatial features of EEG signals. To enhance the robustness of the model, dropout layers are introduced before each parallel DSTS branch. The subsequent sections will provide a detailed description of the structure and function of each module.
3.1 Downsampling projector for EEG Signal Preprocessing
The Downsampling projector module within our EEG-based deep learning framework is designed to preprocess Motor Imagery EEG data, preparing it for intricate analysis by subsequent Transformer and Temporal Convolutional Network (TCN) layers. This module adeptly reshapes high-dimensional EEG sequences, characterized by a temporal resolution of 1125 and 22 channels, into a format that is conducive to convolutional processing. The main purpose of this process is to reduce the length of the sequence by passing continuous EEG signals through simple convolutional layers and average pooling layers.
@@ 3.1 Downsampling projector for EEG Signal Preprocessing
The core architecture of the Downsampling projector, as illustrated in Figure 2, comprises three convolutional layers. The first convolutional layer is designed to initiate the feature extraction process without the application of an activation function. In contrast, the second and third convolutional layers are each followed by a batch normalization (BN) layer and an exponential linear unit (ELU) activation layer to stabilize the learning process and introduce non-linear dynamics into the model. The ELU [24] activation function is defined as:
@@
Figure 2: Architecture of the Downsampling projector. The figure provides a detailed schematic of the Downsampling projector’s architecture. It includes three convolutional layers, with the second and third layers each followed by a batch normalization (BN) layer and an ELU activation layer. Additionally, two average pooling layers and two dropout layers are incorporated to foster model generalization. Specific parameters, such as the kernel size and stride for the convolutional layers, and the kernel size for the average pooling layers, are also depicted. For example, "Conv 1x16, (64,1)" signifies a convolutional layer transitioning from an input channel depth of 1 to an output channel depth of 16, with a stride of 64 along the width and 1 along the height of the input feature map.
@@ 3.2 Stabilizing the Transformer Layer
In the subsequent modules, we employ a modified Transformer layer [25], which has been adapted with recent technological advancements to enhance training stability and model efficacy. Here, we detail the specific alterations applied to the Transformer architecture.
@@
def transformer_block_pre_norm(x):     x = x + self_attention(norm(x))     x = x + feed_forward(norm(x))     return x
@@
Our approach also incorporates RMSNorm, or Root Mean Square Layer Normalization [26], as the normalization function. RMSNorm diverges from traditional Layer Normalization by normalizing solely the standard deviation of activations, not the mean and standard deviation. It achieves this by dividing the activations by their root mean square, which maintains gradient scale and facilitates the training of deep networks.
@@
To further enhance our model, we have substituted the typically-used ReLU activation with the Swish Gated Linear Unit (SwiGLU) [27]. SwiGLU is defined as the componentwise product of two linear transformations of the input, one of which is Swish-activated:
$\text{SwiGLU}(x,W,V,b,c)=Swish_{\beta}(xW+b)\odot(xV+c)$ (3)
In the equation, $x$ is the input, $W$ and $V$ are weight matrices, Swish [28] is defined as $x\cdot\delta(\beta x)$, where $\delta(z)=(1+exp(-z))^{-1}$ is the sigmoid function and $\beta$ is either a constant or a trainable parameter. SwiGLU’s principal advantages are:
@@ 3.3 Dual-Stream Temporal-Spatial Block
The Dual-Stream Temporal-Spatial Block (DSTS Block) , as shown in Figure3 presents an architecture specifically designed for the analysis of electroencephalogram (EEG) data during Motor Imagery (MI) tasks. This architecture integrates Temporal Convolutional Networks (TCNs) with stable Transformer modules, capitalizing on their complementary strengths to capture the temporal and spatial characteristics inherent in EEG signals.
Figure 3: Architecture of the DTDS Block. The DTDS Block integrates a TCN for local temporal feature extraction with a self-attention block for global spatial context analysis, enabling a detailed examination of EEG signals for MI classification tasks.
@@
The DSTS Block is engineered to leverage the TCN’s proficiency in local feature extraction and the Transformer’s capacity for global context comprehension, thus aiming to provide a comprehensive analysis of EEG data. We also adopt the relative position representations as proposed by Shaw et al. in their seminal work[29]. This dual-stream approach is anticipated to improve the model’s ability to identify patterns relevant to MI tasks by enhancing its analytical complexity.
@@ 3.4 EEG Signal Classification with EEGEncoder
4 Results
In this section, we provide a detailed evaluation of the EEGEncoder model, demonstrating its classification capabilities on the BCI Competition IV 2a dataset[30]. We compare the performance of our model with various established models to underscore its effectiveness in decoding the complex patterns inherent in EEG signals for motor imagery tasks. The subsequent subsections elaborate on the model’s performance metrics, a comparative analysis with other models, and discuss the significance of these results for the progression of brain-computer interface technologies.
4.1 Datasets
In our study, we primarily utilized the BCI Competition IV dataset 2a (BCI-2a) for training and evaluating the EEGEncoder model. The BCI-2a dataset comprises recordings from nine subjects across two sessions, with each subject performing 288 motor imagery trials. The trials were recorded using 22 EEG electrodes at a sampling rate of 250 Hz. The data underwent bandpass filtering between 0.5 and 100 Hz and a notch filter at 50 Hz to reduce power line interference.
@@ 4.4 Ablation Study
Here, we present a selection of key experiments that were instrumental in assessing the impact of specific modifications. These experiments included removing the transformer component from the DTDS block, using 5 shift windows instead of five dropout branches, varying the number of transformer layers, adjusting the quantity of DTDS branches within the EEGEncoder, and comparing the performance of our modified stable transformer against the Vanilla Transformer. To ensure the statistical significance of our results, we averaged the outcomes across five iterations, each initialized with a different random seed. The summarized results are displayed in Table 3.
@@
shift window 74.43 64.2
Transformer-2 Layer 74.05 64.4
Transformer-4 Layer 74.04 64.2
Transformer-8 Layer 73.78 64.1